The event you wear flats to school today is A and the event you wear sneakers to school today is B. If these events are mutually exclusive events, using P(A) = 0.34, and P(B) = 0.26, what is P(B/A)?
Answer #1

mutually exclusive events cannot occur at the same time.

so P(A and B)=0 for mutually exclusive

P(B/A)=P(A and B)/P(A)

=0/0.34

=0

ANSWER:

P(B/A)=0